# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ 4.4 Detectability of the SGWB anisotropies
Based on the formulae for the correlation functions of the SGWB anisotropies induced by the inflationary STT bispectrum with the non-Bunch-Davies states shown in the above sections, let us discuss the detectability of such inflationary SGWB anisotropies in future experiments. The detectability of the SGWB anisotropies has been discussed for space-based interferometers BBO/DECIGO and Square Kilometre Array (SKA) as a PTA experiment in Ref. [29], where a noise angular spectrum $N_{\ell}$ for each experiment has been computed for the case of a blue-tilted tensor spectrum. In fact, in our paper, we have considered the peaked tensor power spectrum with the non-Bunch-Davies state to explain the recent SGWB data reported by the PTA experiments such as NANOGrav, and it is not exactly the same as the one assumed in Ref. [29]. However, as discussed in the previous section 2.2.1, by setting the model parameters appropriately, our model can reproduce the blue-tilted tensor power spectrum for the relevant frequency range of the PTA experiments. Thus, we employ their result of the noise spectrum when discussing the detectability with future PTA experiments such as SKA. For the scale-invariant case discussed in Sec. 4.3.2, we employ the noise spectrum computed in Ref. [110]. We investigate the detectability of the SGWB anisotropies for the case of the scale-independent and -dependent Bogoliubov coefficients with BBO/DECIGO and SKA, respectively.
@@
Since the power spectra of the SGWB anisotropies induced from the primordial STT bispectrum with the non-Bunch-Davies states decrease as the multipole moment $\ell$ increases as shown in the previous section, and also the amplitudes of the noise spectra rapidly increase at $\ell>10$ [29, 110], in the estimation of the detectability we only consider the $-\ell$ region as $\ell\leq 10$. Following Refs. [29, 110], for $\ell\leq 10$, the typical amplitudes of the noise spectrum are given as $N_{\ell}=\mathcal{O}(10^{-30})$ in the case of the scale-invariant spectrum for BBO/DECIGO and $N_{\ell}=\mathcal{O}(10^{-26})$ in the case of the blue-tilted spectrum for SKA, respectively. The noise used for computing the signal-to-noise ratio (SNR) of the SGWB anisotropies is
$N_{\ell}^{\rm GW}:=\frac{N_{\ell}}{\bar{\Omega}^{2}_{\rm GW}}.$ (4.54)
@@
For the auto-correlation function, we inspect $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ as the SNR, [^5] Exactly speaking, the SNR squared is evaluated by computing [10, 32] $\sum_{\ell}T_{\rm obs}\int{\rm d}f(2\ell+1)\frac{C_{\ell}^{\rm GW}}{N_{\ell}^{\rm GW}},$ (4.55) where $T_{\rm obs}$ is the total time for observations. In the present paper, we use $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ for the purpose of order estimation. that is, the threshold for the detection is roughly estimated as $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}\gtrsim 1$. In our case, $C_{\ell}^{\rm GW}$ is at most of $\mathcal{O}(10^{-4})$ for both scale-independent and peaked Bogoliubov coefficients. For the DECIGO/BBO case, assuming the scale-invariant power spectrum, we have $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-16})$ for $r=10^{-2}$ and hence $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}\ll 1$. For the PTA case, assuming that the amplitude of the SGWB is that of reported signals with recent PTA experiments, that is, $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$ at nHz, we can expect $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}<O(10^{4})$. Thus, if the reported nHz SGWB is sourced by the inflationary GWs with non-Bunch-Davies state, in future PTA experiments such as SKA it could be possible to test the vacuum state during inflation by the observation of the SGWB anisotropies.
As for the scale-invariant case with DECIGO/BBO experiments, let us briefly investigate the detectability by making use of the cross-correlation between GWs and CMB. In this case, as we have discussed, $C^{\rm GW}_{\ell}\ll N^{\rm GW}_{\ell}$, and then the SNR for the cross-correlation can be roughly computed as [29]  [^6] For the case of $N^{\rm GW}_{\ell}\lesssim C^{\rm GW}_{\ell}$, the SNR can be roughly given by [29] ${\rm SNR}\simeq\frac{C^{\rm GW-T}_{\ell}}{\sqrt{(C^{\rm GW}_{\ell}+N^{\rm GW}_{\ell})C^{\rm TT}_{\ell}}}.$ (4.56) For the SKA case (i.e., $N_{\ell}=\mathcal{O}(10^{-26})$), by assuming $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$, the above SNR can reach unity for $f^{ss}_{{\rm NL},0}\geq\mathcal{O}(1)$. Therefore, in light of the upper bound on $f^{ss}_{{\rm NL},0}$, the amplitudes of the SGWB anisotropies for the peaked-spectrum case can be large enough to be tested by both auto- and cross-correlation.
${\rm SNR}\simeq\frac{C_{\ell}^{\rm GW-T}}{\sqrt{N_{\ell}^{\rm GW}C_{\ell}^{\rm TT}}},$ (4.57)
@@
$C^{\rm TT}_{\ell}=\frac{1}{2\ell+1}\sum_{m}\langle a^{T}_{\ell m}a^{T*}_{\ell m}\rangle\simeq\frac{2\pi}{25\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-11}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (4.58)
Here, we have used the approximated form of the transfer function of the temperature fluctuation in Eq. (4.25). Thus, even if the cross-correlation would be used for the DECIGO/BBO experiments, it is hard to detect the SGWB anisotropies induced from the inflationary STT bispectrum with the scale-independent Bogoliubov coefficient. Note that, similarly to the SKA case, one can consider the tensor spectrum peaked at the frequency band targeted by DECIGO/BBO to enhance $\bar{\Omega}_{\rm GW}$ without contradicting CMB experiments by employing the scale-dependent Bogoliubov coefficient. For such a case, we need to estimate the noise spectrum with the peaked tensor power spectrum for the DECIGO/BBO experiments and also investigate the backreaction problem carefully. We will leave the detailed analysis of these points to our future work.
@@ 5 Summary
The SGWB anisotropies have recently attracted a lot of attention as a tool to distinguish the source of SGWBs and a probe of the new physics in the early universe. In this paper, we have investigated the SGWB anisotropies induced by the primordial non-Gaussianity during inflation. Although it has been mentioned that the characteristic primordial non-Gaussianity with non-Bunch-Davies states, which brings the coupling between short- and long-wavelength modes, can induce the spatial anisotropies on SGWB, precise theoretical predictions were not made in previous studies. In this paper, we have explicitly formulated the primordial scalar-tensor-tensor (STT) bispectrum induced during inflation in the context of Horndeski theory with non-Bunch-Davies states and investigated the STT bispectrum in the squeezed limit, which characterizes the coupling between long-wavelength scalar mode and short-wavelength tensor modes. In the formulation, we carefully clarified the Trans-Planckian problem for the short-wavelength tensor modes in the non-Bunch-Davies state and then derived the condition to avoid it. In the CMB observations, due to the relatively small hierarchy between the longest and shortest modes which are in the observational range (roughly estimated as $k_{\rm S}/k_{\rm L}=O(10^{3})$), the aforementioned Trans-Planckian problem is not so serious. On the other hand, for the SGWB anisotropies, a large hierarchy between long- and short-wavelength modes (roughly estimated as $k_{\rm S}/k_{\rm L}=O(10^{8})$ (for PTA experiments), and $O(10^{16})$ (for DECIGO/BBO experiments)) is required, and then close attention must be paid to this problem. We found that by avoiding the Trans-Planckian problem the strong scale-dependent enhancement of the STT bispectrum in the squeezed limit in the case with the non-Bunch-Davies states, which was considered in the previous works, can not be expected. Instead, another scale dependence appears in the STT bispectrum which also depends on the non-Bunch-Davies excitation time.
@@
To obtain the theoretical predictions for the SGWB anisotropies, with future GW observations in mind, we assumed the two configurations of the Bogoliubov coefficients: one is for the highly blue-tilted tensor spectrum, and the other is for the almost scale-invariant spectrum. Using those configurations, we have computed the self-correlation function of the SGWB anisotropies and the cross-correlation function of those with the CMB temperature fluctuations. Then, we evaluated the possible maximum values of the amplitudes of those correlation functions in light of the backreaction constraint and the perturbativity condition. As a result, we have found that both anisotropies can dominate those originating from the late-time metric perturbations, so-called SW effects, and shown that the anisotropies with the maximum amplitudes can potentially be tested with PTA experiments. We have also found that the SNR for the scale-invariant case is much smaller than unity. However, since how large the SNR can be depends on the magnitude of $\bar{\Omega}_{\rm GW}$, those anisotropies could be tested with DECIGO/BBO as well if a relatively large tensor power spectrum compared to the scale-invariant case is detected at the deci-Hz frequency band. Here, our analysis of the detectability is based on order estimations, and hence it would be important to compute the noise spectrum explicitly and discuss the detectability in detail. We will leave it to our future work.
@@ Acknowledgements
We thank Chunshan Lin and Giorgio Orlando for fruitful discussions. This work is supported by the grant No. UMO-2021/42/E/ST9/00260 from the National Science Centre, Poland (S.A.), MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe”, No. JP21H05189 (S.A), and JSPS KAKENHI Grants No. JP20K03968 (S.Y.), JP23H00108 (S.Y.) and JP24K00627 (S.Y.).
@@ A.1 Quadratic actions
By substituting the perturbed metric Eq. (2.2) into the action Eq. (2.1) and expanding up to the quadratic order, the quadratic action has been obtained as [82]
@@
$\quad\quad\quad\quad\quad\quad\ +\frac{\mathcal{G}_{T}}{8}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr{]},$ (A.1)
@@
$\quad+2H^{3}\dot{\phi}\left(15XG_{5X}+13X^{2}G_{5XX}+2X^{3}G_{5XXX}\right)-6H^{2}X(6G_{5\phi}+9XG_{5\phi{X}}+2X^{2}G_{5\phi{X}X}).$ (A.3)
@@
$\alpha$ $=\frac{\mathcal{G}_{T}}{\Theta}\dot{\zeta},$ (A.4)
$\beta$ $=\frac{1}{a\mathcal{G}_{T}}\biggl{(}a^{3}\mathcal{G}_{S}\psi-\frac{a\mathcal{G}_{T}^{2}}{\Theta}\zeta\biggr{)},$ (A.5)
where $\psi:=\partial^{-2}\dot{\zeta}$. Eliminating $\alpha$ and $\beta$ in the quadratic actions by using the solutions of the constraint equations, one obtains the quadratic actions in Eq. (2.5) and Eq. (2.6) where the coefficients are of the form,
$\mathcal{G}_{T}$ $:=2\left[G_{4}-2XG_{4X}-X(H\dot{\phi}G_{5X}-G_{5\phi})\right],$ (A.6)
$\mathcal{F}_{T}$ $:=2\left[G_{4}-X(\ddot{\phi}G_{5X}+G_{5\phi})\right],$ (A.7)
$\mathcal{G}_{S}$ $:=\mathcal{G}_{T}\biggl{(}\frac{\Sigma\mathcal{G}_{T}}{\Theta^{2}}+3\biggr{)},$ (A.8)
$\mathcal{F}_{S}$ $:=\frac{1}{a}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{a\mathcal{G}_{T}^{2}}{\Theta}\biggr{)}-\mathcal{F}_{T}.$ (A.9)
A.2 Cubic action from scalar-tensor-tensor interactions
Similarly to the quadratic actions, after expanding the Horndeski action with respect to metric perturbations up to the cubic order, the scalar-tensor-tensor type cubic action can be obtained as [101]

### v3

@@ 4.4 Detectability of the SGWB anisotropies
Based on the formulae shown in the above sections for the correlation functions of the SGWB anisotropies induced by the inflationary STT bispectrum with the non-Bunch-Davies states, let us discuss the detectability of such inflationary SGWB anisotropies in future experiments. The detectability of the SGWB anisotropies has been discussed for the space-based interferometers BBO/DECIGO and Square Kilometre Array (SKA) as a PTA experiment in Ref. Dimastrogiovanni:2021mfs , where a noise angular spectrum $N_{\ell}$ for each experiment has been computed for the case of a blue-tilted tensor spectrum. In fact, in our paper, we have considered the peaked tensor power spectrum with the non-Bunch-Davies state to explain the recent SGWB data reported by the PTA experiments such as NANOGrav, and it is not exactly the same as the one assumed in Ref. Dimastrogiovanni:2021mfs . However, as discussed in the previous section 2.2.1, by setting the model parameters appropriately, our model can reproduce the blue-tilted tensor power spectrum for the relevant frequency range of the PTA experiments. Thus, we employ their result of the noise spectrum when discussing the detectability with future PTA experiments such as SKA. For the scale-invariant case discussed in Sec. 4.3.2, we employ the noise spectrum computed in Ref. Ricciardone:2021kel . We investigate the detectability of the SGWB anisotropies for the case of the scale-independent and -dependent Bogoliubov coefficients with BBO/DECIGO and SKA, respectively.
@@
Since the power spectra of the SGWB anisotropies induced from the primordial STT bispectrum with the non-Bunch-Davies states decrease as the multipole moment $\ell$ increases as shown in the previous section, and also the amplitudes of the noise spectra rapidly increase at $\ell>10$ Dimastrogiovanni:2021mfs ; Ricciardone:2021kel , in the estimation of the detectability we only consider the low-$\ell$ region as $\ell\leq 10$. Following Refs. Dimastrogiovanni:2021mfs ; Ricciardone:2021kel , for $\ell\leq 10$, the typical amplitudes of the noise spectrum are given as $N_{\ell}=\mathcal{O}(10^{-30})$ in the case of the scale-invariant spectrum for BBO/DECIGO and $N_{\ell}=\mathcal{O}(10^{-26})$ in the case of the blue-tilted spectrum for SKA, respectively. The noise used for computing the signal-to-noise ratio (SNR) of the SGWB anisotropies is
$N_{\ell}^{\rm GW}:=\frac{N_{\ell}}{\bar{\Omega}^{2}_{\rm GW}}.$ (135)
@@
For the auto-correlation function, we inspect $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ as the SNR, [^5] Exactly speaking, the SNR squared is evaluated by computing LISACosmologyWorkingGroup:2022jok ; Orlando:2022rih : $\sum_{\ell}T_{\rm obs}\int{\rm d}f(2\ell+1)\frac{C_{\ell}^{\rm GW}}{N_{\ell}^{\rm GW}},$ (136) where $T_{\rm obs}$ is the total time for observations. In the present paper, we use $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}$ for the purpose of order estimation. i.e. the threshold for detection is roughly estimated as $C^{\rm GW}_{\ell}/N_{\ell}^{\rm GW}\gtrsim 1$. In our case where the cutoff scale is given by Eq. (35), $C_{\ell}^{\rm GW}$ can increase up to $\mathcal{O}(10^{4})$ and $\mathcal{O}(10^{-2})$ for the scale-independent and peaked Bogoliubov coefficients, respectively. For the DECIGO/BBO case, assuming the scale-invariant power spectrum, we have $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-16})$ for $r=10^{-2}$ and hence $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}\leq\mathcal{O}(10^{2})$. Therefore, if DECIGO/BBO would detect SGWBs, it could be possible to test the vacuum state during inflation by the anisotropies of those SGWBs. Note that the maximum possible value of the SNR depends on the cutoff scale, as we have discussed. For the PTA case, assuming that the amplitude of the SGWBs is that of reported signals with recent PTA experiments, i.e. $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$ at nHz, we can expect $C_{\ell}^{\rm GW}/N_{\ell}^{\rm GW}\leq O(10^{6})$, and hence, if the reported nHz SGWB is sourced by the inflationary GWs with non-Bunch-Davies states, in future PTA experiments such as SKA it could be possible to test the vacuum state by the SGWB anisotropies, similarly to DECIGO/BBO.
In the following, let us briefly investigate the detectability by making use of the cross-correlation between GWs and CMB. The SNR for the cross-correlation can be roughly computed as Dimastrogiovanni:2021mfs :
${\rm SNR}\simeq\frac{C^{\rm GW-T}_{\ell}}{\sqrt{(C^{\rm GW}_{\ell}+N^{\rm GW}_{\ell})C^{\rm TT}_{\ell}}},$ (137)
@@
$C^{\rm TT}_{\ell}=\frac{1}{2\ell+1}\sum_{m}\langle a^{T}_{\ell m}a^{T*}_{\ell m}\rangle\simeq\frac{2\pi}{25\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-11}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (138)
Here, we have used the approximated form of the transfer function of the temperature fluctuation in Eq. (107). We then find that the SNR can be $\mathcal{O}(1)$ for $f^{ss}_{{\rm NL},0}\geq\mathcal{O}(10^{6})$. For the SKA case (i.e. $N_{\ell}=\mathcal{O}(10^{-26})$), by assuming $\bar{\Omega}_{\rm GW}=\mathcal{O}(10^{-9})$, the above SNR can be $\mathcal{O}(1)$ for $f^{ss}_{{\rm NL},0}\geq\mathcal{O}(1)$. Thus, it could be possible to test the vacuum state during inflation by the cross-correlation as well as the auto-correlation. Note that, similarly to the SKA case, one can consider the tensor spectrum peaked at the frequency band targeted by DECIGO/BBO to enhance $\bar{\Omega}_{\rm GW}$ (i.e. suppress $N^{\rm GW}_{\ell}$) without contradicting CMB experiments by employing the scale-dependent Bogoliubov coefficient. For such a case, we need to estimate the noise spectrum with the peaked tensor power spectrum for the DECIGO/BBO experiments and also investigate the backreaction problem carefully. We will leave the detailed analysis of these points to our future work.
Before closing this section, let us mention the detectability by DECIGO/BBO for higher cutoff scales. In the present paper, we chose the lowest cutoff scale to obtain the most conservative constraint on the excitation time, but the backreaction constraint, the constraint on the non-linearity parameter, and the maximum value of the SNR are all sensitive to the choice of the cutoff scale. Indeed, the maximum possible value of the SNR for a higher cutoff scale can be smaller than unity, and thus it is important to investigate a strong coupling scale for each inflation model to discuss the detectability of the SGWB anisotropies predicted from each one.
@@ 5 Summary
The SGWB anisotropies have recently attracted a lot of attention as a tool to distinguish the source of SGWBs and a probe of the new physics in the early universe. In this paper, we have investigated the SGWB anisotropies induced by the primordial non-Gaussianity during inflation. Although it has been mentioned that the characteristic primordial non-Gaussianity with non-Bunch-Davies states, which brings the coupling between short- and long-wavelength modes, can induce the spatial anisotropies on SGWB, precise theoretical predictions were not made in previous studies. In this paper, we have explicitly formulated the primordial STT bispectrum induced during inflation in the context of Horndeski theory with non-Bunch-Davies states and investigated the STT bispectrum in the squeezed limit, which characterizes the coupling between long-wavelength scalar modes and short-wavelength tensor modes. In the formulation, we carefully clarified the Trans-Planckian and strong coupling problems for the short-wavelength tensor modes in the non-Bunch-Davies state and then derived the condition to avoid them. In particular, those problems were found to be crucial for the modes satisfying $k_{S}/k_{L}>\mathcal{O}(10)$. For the SGWB anisotropies, a large hierarchy between long- and short-wavelength modes (roughly estimated as $k_{\rm S}/k_{\rm L}=O(10^{8})$ (for PTA experiments) and $O(10^{16})$ (for DECIGO/BBO experiments)) is required, and thus close attention must be paid to them. Here, even in the CMB observations with the relatively small hierarchy between the longest and shortest modes in the observational range (roughly estimated as $k_{\rm S}/k_{\rm L}=\mathcal{O}(10^{3})$), such issues are crucial. Therefore, depending on the theory and the model that determine the strong coupling scale, one must pay attention to them. We then found that by avoiding both the Trans-Planckian and strong coupling problems the strong scale-dependent enhancement of the STT bispectrum in the squeezed limit in the case with the non-Bunch-Davies states, which was considered in the previous works, cannot be expected. Instead, another scale dependence appears in the STT bispectrum which also depends on the non-Bunch-Davies excitation time.
@@
To obtain the theoretical predictions for the SGWB anisotropies, with future GW observations in mind, we assumed the two configurations of the Bogoliubov coefficients: one for the highly blue-tilted tensor spectrum, and the other for the almost scale-invariant spectrum. Using those configurations, we computed the self-correlation function of the SGWB anisotropies and the cross-correlation function of those with the CMB temperature fluctuations. Then, we evaluated the maximum possible values of the amplitudes of those correlation functions in light of the backreaction constraint and the perturbativity condition. As a result, we have found that both anisotropies can dominate those originating from the late-time metric perturbations, so-called SW effects, and shown that the anisotropies with the maximum amplitudes can potentially be tested with the DECIGO/BBO and PTA experiments. Here, ${\rm SNR}\geq 1$ for DECIGO/BBO requires a relatively large non-linearity parameter such as $\mathcal{O}(10^{6})$, compared to the PTA experiments case which requires the non-linearity parameter to be $\geq\mathcal{O}(1)$ for ${\rm SNR}\geq 1$. Such a large value of the non-linearity parameter might imply the necessity of taking into account higher-point correlation functions such as a trispectrum in evaluating the non-Gaussianities. Since how large the SNR can be depends on the magnitude of $\bar{\Omega}_{\rm GW}$, the anisotropies with sufficiently large amplitudes detected by DECIGO/BBO would be obtained with a smaller value of the non-linearity parameter if a relatively large tensor power spectrum compared to the scale-invariant case is obtained. Therefore, it would be worth applying the peaked Bogoliubov coefficient to SGWBs at deci-Hz. Also, our analysis of the detectability is based on order estimations, and hence it would be important to compute the noise spectrum explicitly and discuss the detectability in detail. In addition, it would be interesting to discuss the ability to distinguish between models with future experiments based on the SGWB anisotropies. We will leave those to our future work.
@@ Acknowledgements
We thank Chunshan Lin and Giorgio Orlando for fruitful discussions. We thank the anonymous referee for useful comments. This work is supported by the grant No. UMO-2021/42/E/ST9/00260 from the National Science Centre, Poland (S.A.), MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe”, No. JP21H05189 (S.A), and JSPS KAKENHI Grants No. JP20K03968 (S.Y.), JP23H00108 (S.Y.) and JP24K00627 (S.Y.).
@@ A.1 Quadratic actions
By substituting the perturbed metric Eq. (2) into the action Eq. (1) and expanding up to the quadratic order, the quadratic action has been obtained as Kobayashi:2011nu :
@@
$\quad\quad\quad\quad\quad\quad\ +\frac{\mathcal{G}_{T}}{8}\dot{h}_{ij}^{2}-\frac{\mathcal{F}_{T}}{a^{2}}(\partial_{k}h_{ij})^{2}\biggr],$ (139)
@@
$\quad+2H^{3}\dot{\phi}\left(15XG_{5X}+13X^{2}G_{5XX}+2X^{3}G_{5XXX}\right)-6H^{2}X(6G_{5\phi}+9XG_{5\phi{X}}+2X^{2}G_{5\phi{X}X}).$ (141)
@@
$\alpha$ $=\frac{\mathcal{G}_{T}}{\Theta}\dot{\zeta},$ (142)
$\beta$ $=\frac{1}{a\mathcal{G}_{T}}\biggl(a^{3}\mathcal{G}_{S}\psi-\frac{a\mathcal{G}_{T}^{2}}{\Theta}\zeta\biggr),$ (143)
where $\psi:=\partial^{-2}\dot{\zeta}$. Eliminating $\alpha$ and $\beta$ in the quadratic actions by using the solutions of the constraint equations, one obtains the quadratic actions in Eq. (5) and Eq. (6) where the coefficients are of the form,
$\mathcal{G}_{T}$ $:=2\left[G_{4}-2XG_{4X}-X(H\dot{\phi}G_{5X}-G_{5\phi})\right],$ (144)
$\mathcal{F}_{T}$ $:=2\left[G_{4}-X(\ddot{\phi}G_{5X}+G_{5\phi})\right],$ (145)
$\mathcal{G}_{S}$ $:=\mathcal{G}_{T}\biggl(\frac{\Sigma\mathcal{G}_{T}}{\Theta^{2}}+3\biggr),$ (146)
$\mathcal{F}_{S}$ $:=\frac{1}{a}\frac{{\rm d}}{{\rm d}t}\biggl(\frac{a\mathcal{G}_{T}^{2}}{\Theta}\biggr)-\mathcal{F}_{T}.$ (147)
A.2 Cubic action from STT interactions
Similarly to the quadratic actions, after expanding the Horndeski action with respect to metric perturbations up to the cubic order, the STT-type cubic action can be obtained as Gao:2012ib :